### v1

@@ Abstract
Recent advances in multimodal large language models (MLLMs) have catalyzed transformative progress in affective computing, enabling models to exhibit emergent emotional intelligence. Despite substantial methodological progress, current emotional benchmarks remain limited, as it is still unknown: (a) the generalization abilities of MLLMs across distinct scenarios, and (b) their reasoning capabilities to identify the triggering factors behind emotional states. To bridge these gaps, we present MME-Emotion, a systematic benchmark that assesses both emotional understanding and reasoning capabilities of MLLMs, enjoying scalable capacity, diverse settings, and unified protocols. As the largest emotional intelligence benchmark for MLLMs, MME-Emotion contains over 6,000 curated video clips with task-specific questioning-answering (QA) pairs, spanning broad scenarios to formulate eight emotional tasks. It further incorporates a holistic evaluation suite with hybrid metrics for emotion recognition and reasoning, analyzed through a multi-agent system framework. Through a rigorous evaluation of 20 advanced MLLMs, we uncover both their strengths and limitations, yielding several key insights: ❶ Current MLLMs exhibit unsatisfactory emotional intelligence, with the best-performing model achieving only $39.3\%$ recognition score and $56.0\%$ Chain-of-Thought (CoT) score on our benchmark. ❷ Generalist models (e.g., Gemini-2.5-Pro) derive emotional intelligence from generalized multimodal understanding capabilities, while specialist models (e.g., R1-Omni) can achieve comparable performance through domain-specific post-training adaptation. By introducing MME-Emotion, we hope that it can serve as a foundation for advancing MLLMs’ emotional intelligence in the future. Project Page: https://mme-emotion.github.io/.
@@ 1 Introduction
Affective computing Picard (2000); Tao & Tan (2005) is a fascinating interdisciplinary field seeking to bridge the gap between human emotional intelligence and machine capabilities, with significant applications in education Wu et al. (2016); Yadegaridehkordi et al. (2019), healthcare Yannakakis (2018); Liu et al. (2024b), and human-robot interaction Filippini et al. (2020); Gervasi et al. (2023). Catalyzed by emergent multimodal large language models (MLLMs) Wu et al. (2023); Zhang et al. (2024b), research focus has increasingly shifted from exploring emotion recognition capability Kosti et al. (2017) to investigating the interpretability of clues behind emotional states Lian et al. (2024c).
@@ 2.1 Emotional Intelligence in Multimodal Large Language Models
The success of MLLMs Liu et al. (2023a); Zhu et al. (2023) has advanced affective computing, prompting growing interest in equipping these models with emotional intelligence Cheng et al. (2024a); Zhao et al. (2025b). On the one hand, general MLLMs (such as the Gemini Team et al. (2023), GPT Achiam et al. (2023), and Qwen-VL Bai et al. (2023) series) have demonstrated strong multimodal understanding capabilities. Thanks to their powerful LLM backbones and extensive pretraining on diverse datasets, emotional intelligence emerges as a byproduct of these general models. On the other hand, MLLMs with relatively small parameter sizes (ranging from 0.5B to 7B) have not yet shown strong general multimodal understanding capabilities, primarily due to limitations in data scale and model capacity. However, with post-training adaptation strategies, it is still possible to elicit strong performance on emotion-related tasks. For example, Emotion-LLaMA Cheng et al. (2024a) employs emotion-specific encoders to project video, audio, and text modalities into a unified LLM space, followed by instruction tuning to activate emotional intelligence. R1-Omni Zhao et al. (2025a) enhances emotional reasoning capabilities in MLLMs through reinforcement learning with verified feedback (RLVR). AffectGPT Lian et al. (2024c), by training on a large-scale emotional dataset and introducing a pre-fusion projector to integrate multimodal emotional signals, builds a powerful emotion-specialized model. However, these models are often evaluated only on specific tasks and limited scenarios, leaving it unclear how they perform under open and fair evaluation settings.
@@ 3.2 Evaluation Suite
Next, we provide a detailed introduction to the evaluation metrics employed in this paper. During the earlier step extraction process from MLLM answers, we consistently place the task-specific prediction in the final step, while treating all other steps as the reasoning process. To compute the recognition score, we compare the final prediction step against the ground-truth emotion labels. For tasks involving a single emotion label, we adopt standard accuracy as the recognition score, following conventional evaluation practices Li & Deng (2020); Tzirakis et al. (2017). For tasks involving multiple emotion labels, the recognition score is defined as the ratio between the number of correctly predicted emotions and the total number of ground-truth emotions. We then derive the reasoning score from the extracted reasoning steps. Each step is treated as a binary classification problem, where the judge agent determines whether the reasoning is correct or not. To eliminate potential bias caused by varying step lengths, we compute the reasoning score as the average correctness across all reasoning steps. Finally, we obtain the CoT score by taking a weighted combination of the recognition score and the reasoning score, providing a holistic measure of both recognition accuracy and reasoning quality. The CoT score is defined as:
@@
Colsed-source MLLMs
@@ 4.3 Observations and Insights
On the one hand, generalist MLLMs, such as the Gemini and GPT series, are equipped with large model capacities and trained on massive datasets. As a result, they exhibit strong general understanding capabilities and perform well across various tasks Li et al. (2024); Fu et al. (2023); Zhang et al. (2024c); Fu et al. (2025); Jiang et al. (2025). Although these models have not been explicitly adapted to emotion-specific domains Fu et al. (2024), our experimental results show that they can generalize from general intelligence to emotional intelligence, illustrating one viable path forward. On the other hand, specialist models, such as AffectGPT and R1-Omni, may not match generalist models in broad understanding domains due to the limitation of scaling laws Kaplan et al. (2020); Zhang et al. (2024a). However, by constructing large-scale emotional data and applying post-training adaptation, they can achieve comparable performance against generalist models. These findings demonstrate that there is no one-size-fits-all solution: emotional intelligence can be bootstrapped through a variety of strategies, opening the door to a broad range of future research directions.
@@ 5 Conclusion
Appendix A Limitations and Social Impact
@@
Appendix B More Related Works
B.1 Datasets and Benchmarks for Multimodal Large Language Models
With the rise of large language models (LLMs) and the growing popularity of multimodal large language models (MLLMs), an increasing number of studies have focused on evaluating MLLM performance across a wide range of domains. By collecting data and employing human annotations, MME Fu et al. (2023) evaluates the general understanding capabilities of MLLMs across 14 subtasks, ranging from perception tasks to cognition tasks. Video-MME Fu et al. (2025) assesses the video analysis capabilities of MLLMs using a diverse and high-quality collection of video data. MME-RealWorld Zhang et al. (2024c) aggregates a large set of high-resolution images, combined with manual annotations, to construct a suite of practical and challenging tasks, including optical character recognition (OCR) in the wild, remote sensing, and autonomous driving. FinMME Luo et al. (2025) leverages annotations from 20 domain experts to comprehensively evaluate MLLMs’ understanding capabilities in the financial domain. These benchmarks have significantly accelerated progress in MLLM development, enabling improved performance in both general-purpose and domain-specific scenarios, ultimately contributing to a wide array of real-world applications.
Appendix C Implementation Details
C.1 Evaluated MLLMs
@@
C.2 Prompt for Evaluated MLLMs
@@
C.3 Prompt for Extracting Audio Clues
@@
C.4 Prompt for Extracting Key Steps
@@
C.5 Prompt for Rating Performance
@@
Appendix D Additional Experimental Results
D.1 Additional Experimental Results across Tasks
@@
D.2 Additional Experimental Results across Categories
@@
D.3 Additional Experimental Results across Metrics
@@
D.4 Additional Case Study
@@
Colsed-source MLLMs
@@
Colsed-source MLLMs
@@
Colsed-source MLLMs
@@
Colsed-source MLLMs
@@
Colsed-source MLLMs
@@
Colsed-source MLLMs
@@
Colsed-source MLLMs
@@
Colsed-source MLLMs

### v2

@@ Abstract
Recent advances in multimodal large language models (MLLMs) have catalyzed transformative progress in affective computing, enabling models to exhibit emergent emotional intelligence. Despite substantial methodological progress, current emotional benchmarks remain limited, as it is still unknown: (a) the generalization abilities of MLLMs across distinct scenarios, and (b) their reasoning capabilities to identify the triggering factors behind emotional states. To bridge these gaps, we present MME-Emotion, a systematic benchmark that assesses both emotional understanding and reasoning capabilities of MLLMs, enjoying scalable capacity, diverse settings, and unified protocols. As the largest emotional intelligence benchmark for MLLMs, MME-Emotion contains over 6,000 curated video clips with task-specific questioning-answering (QA) pairs, spanning broad scenarios to formulate eight emotional tasks. It further incorporates a holistic evaluation suite with hybrid metrics for emotion recognition and reasoning, analyzed through a multi-agent system framework. Through a rigorous evaluation of 20 advanced MLLMs, we uncover both their strengths and limitations, yielding several key insights: ❶ Current MLLMs exhibit unsatisfactory emotional intelligence, with the best-performing model achieving only $39.3\%$ recognition score and $56.0\%$ Chain-of-Thought (CoT) score on our benchmark. ❷ Generalist models (e.g., Gemini-2.5-Pro and GPT-4o) derive emotional intelligence from generalized multimodal understanding capabilities, while specialist models (e.g., R1-Omni and AffectGPT) can achieve comparable performance through domain-specific post-training adaptation. By introducing MME-Emotion, we hope that it can serve as a foundation for advancing the emotional intelligence of MLLMs in the future. Project Page: https://mme-emotion.github.io/.
@@ 1 Introduction
Affective computing Picard (2000); Tao and Tan (2005); Zhang et al. (2024c; 2025a); Mao et al. (2025) is a fascinating interdisciplinary field seeking to bridge the gap between human emotional intelligence and machine capabilities, with significant applications in education Wu et al. (2016); Yadegaridehkordi et al. (2019), healthcare Yannakakis (2018); Liu et al. (2024b); Zhang et al. (2025b); Zhu et al. (2025), and human-robot interaction Filippini et al. (2020); Gervasi et al. (2023). Catalyzed by emergent multimodal large language models (MLLMs) Wu et al. (2023); Zhang et al. (2024b), research focus has increasingly shifted from exploring emotion recognition capability Kosti et al. (2017) to investigating the interpretability of clues behind emotional states Lian et al. (2024c).
@@ 2.1 Emotional Intelligence in Multimodal Large Language Models
The success of MLLMs Liu et al. (2023a); Zhu et al. (2023); Luo et al. (2025b); Wang et al. (2025) has advanced affective computing, prompting growing interest in equipping these models with emotional intelligence Cheng et al. (2024a); Zhao et al. (2025b); Lian et al. (2025a); Zhu et al. ; Lian et al. (2025b). On the one hand, general MLLMs (such as the Gemini Team et al. (2023), GPT Achiam et al. (2023), and Qwen-VL Bai et al. (2023) series) have demonstrated strong multimodal understanding capabilities. Thanks to their powerful LLM backbones and extensive pretraining on diverse datasets, emotional intelligence emerges as a byproduct of these general models. On the other hand, MLLMs with relatively small parameter sizes (ranging from 0.5B to 7B) have not yet shown strong general multimodal understanding capabilities, primarily due to limitations in data scale and model capacity. However, with post-training adaptation strategies, it is still possible to elicit strong performance on emotion-related tasks. For example, Emotion-LLaMA Cheng et al. (2024a) employs emotion-specific encoders to project video, audio, and text modalities into a unified LLM space, followed by instruction tuning to activate emotional intelligence. R1-Omni Zhao et al. (2025a) enhances emotional reasoning capabilities in MLLMs through reinforcement learning with verified feedback (RLVR). AffectGPT Lian et al. (2024c), by training on a large-scale emotional dataset and introducing a pre-fusion projector to integrate multimodal emotional signals, builds a powerful emotion-specialized model. However, these models are often evaluated only on specific tasks and limited scenarios, leaving it unclear how they perform under open and fair evaluation settings.
@@ 3.2 Evaluation Suite
Next, we provide a detailed introduction to the evaluation metrics employed in this paper. During the earlier step extraction process from MLLM answers, we consistently place the task-specific prediction in the final step, while treating all other steps as the reasoning process. To compute the recognition score, we compare the final prediction step against the ground-truth emotion labels. For tasks involving a single emotion label, we adopt standard accuracy as the recognition score, following conventional evaluation practices Li and Deng (2020); Tzirakis et al. (2017). For tasks involving multiple emotion labels, the recognition score is defined as the ratio between the number of correctly predicted emotions and the total number of ground-truth emotions. We then derive the reasoning score from the extracted reasoning steps. Each step is treated as a binary classification problem, where the judge agent determines whether the reasoning is correct or not. To eliminate potential bias caused by varying step lengths, we compute the reasoning score as the average correctness across all reasoning steps. Finally, we obtain the CoT score by taking a weighted combination of the recognition score and the reasoning score, providing a holistic measure of both recognition accuracy and reasoning quality. The CoT score is defined as:
@@
Closed-source MLLMs
@@ 4.3 Observations and Insights
On the one hand, generalist MLLMs, such as the Gemini and GPT series, are equipped with large model capacities and trained on massive datasets. As a result, they exhibit strong general understanding capabilities and perform well across various tasks Li et al. (2024); Fu et al. (2023); Zhang et al. (2024d); Fu et al. (2025); Jiang et al. (2025). Although these models have not been explicitly adapted to emotion-specific domains Fu et al. (2024), our experimental results show that they can generalize from general intelligence to emotional intelligence, illustrating one viable path forward. On the other hand, specialist models, such as AffectGPT and R1-Omni, may not match generalist models in broad understanding domains due to the limitation of scaling laws Kaplan et al. (2020); Zhang et al. (2024a). However, by constructing large-scale emotional data and applying post-training adaptation, they can achieve comparable performance against generalist models. These findings demonstrate that there is no one-size-fits-all solution: emotional intelligence can be bootstrapped through a variety of strategies, opening the door to a broad range of future research directions.
@@ 5 Conclusion
Acknowledgment
The work described in this paper was supported in part by the Research Grants Council of the Hong Kong Special Administrative Region, China, under Project T45-401/22-N and Project CUHK 14202125.
Appendix A The Use of Large Language Models (LLMs)
LLMs were not involved in the research ideation or the writing of this paper.
Appendix B Limitations and Social Impact
@@
Appendix C More Related Works
C.1 Datasets and Benchmarks for Multimodal Large Language Models
With the rise of large language models (LLMs) and the growing popularity of multimodal large language models (MLLMs), an increasing number of studies have focused on evaluating MLLM performance across a wide range of domains. By collecting data and employing human annotations, MME Fu et al. (2023) evaluates the general understanding capabilities of MLLMs across 14 subtasks, ranging from perception tasks to cognition tasks. Video-MME Fu et al. (2025) assesses the video analysis capabilities of MLLMs using a diverse and high-quality collection of video data. MME-RealWorld Zhang et al. (2024d) aggregates a large set of high-resolution images, combined with manual annotations, to construct a suite of practical and challenging tasks, including optical character recognition (OCR) in the wild, remote sensing, and autonomous driving. FinMME Luo et al. (2025a) leverages annotations from 20 domain experts to comprehensively evaluate MLLMs’ understanding capabilities in the financial domain. These benchmarks have significantly accelerated progress in MLLM development, enabling improved performance in both general-purpose and domain-specific scenarios, ultimately contributing to a wide array of real-world applications.
Appendix D Implementation Details
D.1 Evaluated MLLMs
@@
D.2 Prompt for Evaluated MLLMs
@@
D.3 Prompt for Extracting Audio Clues
@@
D.4 Prompt for Extracting Key Steps
@@
D.5 Prompt for Rating Performance
@@
Appendix E Additional Experimental Results
E.1 Additional Experimental Results across Tasks
@@
E.2 Additional Experimental Results across Categories
@@
E.3 Additional Experimental Results across Metrics
@@
E.4 Additional Case Study
@@
Closed-source MLLMs
@@
Closed-source MLLMs
@@
Closed-source MLLMs
@@
Closed-source MLLMs
@@
Closed-source MLLMs
@@
Closed-source MLLMs
@@
Closed-source MLLMs
@@
Closed-source MLLMs